# GraphSnapShot: Graph Machine Learning Acceleration with Fast Storage and Retrieval

### v2

@@
GraphSnapShot: Graph Machine Learning Acceleration with Fast Storage and Retrieval
@@ 3 Motivation
In machine learning, masks are powerful tools for reducing computation and enhancing model efficiency. In attention masks of long sequences, models selectively attend to relevant tokens within sequences, bypassing unnecessary computations and enabling more focused processing. Masks can also find utility in regularization techniques like dropout, where they aid in randomly deactivating units during training, thus preventing overfitting and promoting generalization. Additionally, in tasks such as instance segmentation in computer vision, masks efficiently delineate object boundaries, streamlining the computational load by restricting processing to specific regions of interest. Through these mechanisms, masks can reduce computational resources and contribute to the scalability of machine learning models. Notably, masked language modeling [STQ+19] [WGZC23], and masked vision learning [LFH+23] showcase the significance of masks.
@@ 4 Model Construction
Traditional graph learning methods struggle with extensive memory usage and slow computation. Node-wise sampling methods, including well-known ones like GraphSAGE [HYL18], are not optimized for multi-hop scenarios due to the computational resources usage that grows exponentially with each additional hop— a phenomenon known as the ”neighbor explosion.” This makes them unsuitable for large-scale graphs where the capture of extended neighborhoods is crucial. Layer-wise sampling, such as in FastGCN[CMX18], although designed to mitigate this explosion by sampling nodes per layer, often loses valuable connectivity information, leading to a sparse and inaccurate representation of the graph’s multi-hop structure. Furthermore, subgraph sampling methods such as Cluster-GCN [CLS+19] focus on localized computations and overlook vital inter-subgraph connections, failing to capture the broader topology necessary for a comprehensive multi-hop domain analysis.
@@
GraphSnapShot is designed to solve the inefficacy for current graph masking methods on large graphs. Typical distributed graph processing System such as Marius [MWX+21], will re-sample all the local structure each time, and load corresponding node embeddings from disk, which is time-consuming. GraphSnapShot proposing a new method for quick local structure retrieval by integrating the benefits of both static and dynamic processing. In the preprocessing phase, a static graph snapshot is sampled and stored in memory from the disk. Then, in the computation phase, the algorithm re-samples the graph snapshot, using a mix of data partially from memory and partially from disk. Concurrently, a smaller, dynamic snapshot is sampled from the disk. This dynamic snapshot is crucial as it is used to swap with the portion of the graph stored in memory. This process ensures the computation can always learn ”fresh” snapshot without overwhelming the memory capacity. GraphSnapShot allows the algorithm to continuously update and adapt to changes in the graph while maintaining a manageable memory footprint, making it an effective tool for processing large, dynamic graph data.
@@ 5.2 OTF
The ”On The Fly” (OTF) method, updates portions of the cache dynamically while leaving other segments of the cache information unchanged. In detailed implementation of OTF, we design 4 different mode of it for real world application. They are (Full Fetch, Partial Fetch)x(Full Refresh, Partial Refresh).
@@ 5.5 FCR
The third method is FCR, which is short for ”fully cache refreshing” in a batch. For FCR, in each batch, the the structure information are sampled from cache for computation, and between each batch, the cache is fully refreshed.
@@ 6 GraphSnapShot Overview
We use the Deep Graph Library (DGL[WZY+20]) framework to implement new samplers in GraphSnapShot, leveraging its efficient handling of graph data structures and operations. Our implementation involves loading the graph, applying the graph splitting based on node degree, and then processing each subgraph with the respective sampling technique. We evaluated the performance of our new models on datasets in ogbn-benchmark[HFZ+21]. The results show that GraphSnapShot can significantly reduce training time and memory usage without compromising the quality of GNN training.
@@ 7.1 Theoretial Analysis of Traditional Disk-Memory Strategy and Our Disk-Cache-Memory Strategy
In the traditional Graph System like Marius[MWX+21], a disk-cache model is deployed, where the graph structure is resampled each time from the disk into the memory for computation. In our current GraphSnapShot model, we deploy a disk-cache-memory model where K-V pair for reused graph structure are stored in the cache to promote the system performance.
@@ 7.6 Experimental Analysis of SSDReS Performance
Training Time and refresh rate: Our experiments indicated the existence of an optimal refresh rate value between 0 and 1, which minimizes training time, and with a little loss in acuracy. We call this point as ”tradeoff” point between static and dynamic resampling. Over-dependence on dynamic resampling (low refresh rate) could increase the computational burden due to frequent updates required for rapidly changing graph structures, while excessive reliance on static resampling (high refresh rate) might cause the model to overly learn from limited portion of the graph structure in memory, deduce its ability to effectively learn from the entire graph structure.

### v3

@@
Abstract
Graph learning has become essential across various domains, yet current sampling techniques face significant challenges when scaling to dynamic or large-scale networks. Popular methods, such as GraphSAGE and Cluster-GCN, encounter issues like excessive memory usage, neighbor explosion, and limited adaptability to dynamic changes. To address these limitations, we introduce GraphSnapShot, a framework specifically designed for efficient and scalable graph learning through dynamic snapshot caching.
GraphSnapShot incorporates the GraphSDSampler, a key module that dynamically captures, updates, and stores local graph snapshots, optimizing both memory usage and computational efficiency. The framework offers innovative caching strategies, including Fully Cache Refresh (FCR), On-The-Fly (OTF), and Shared Cache Mode, tailored to balance computational demands and cache freshness. OTF methods dynamically update only portions of the graph cache, combining static pre-sampling with real-time dynamic resampling, significantly improving computational speed without compromising accuracy.
We validate the effectiveness of GraphSnapShot on well-known graph learning datasets, including OGBN-arxiv, OGBN-products, OGBN-mag, as well as Citeseer, Cora, and Pubmed. Experimental results show that GraphSnapShot achieves significant reductions in GPU memory usage and computational time compared to traditional samplers, such as those in DGL. Notably, GraphSnapShot maintains competitive accuracy while efficiently handling complex multi-hop neighborhoods in large-scale graphs.
This framework is particularly effective for applications such as social network analysis, bioinformatics, and recommendation systems. Its ability to process evolving graph structures makes it a valuable tool for researchers and practitioners. The code for GraphSnapShot is publicly available at code link111https://github.com/NoakLiu/GraphSnapShot.
@@ 3 Motivation
In machine learning, masks are powerful tools for reducing computation and enhancing model efficiency. In attention masks of long sequences, models selectively attend to relevant tokens within sequences, bypassing unnecessary computations and enabling more focused processing. Masks can also find utility in regularization techniques like dropout, where they aid in randomly deactivating units during training, thus preventing overfitting and promoting generalization. Additionally, in tasks such as instance segmentation in computer vision, masks efficiently delineate object boundaries, streamlining the computational load by restricting processing to specific regions of interest. Through these mechanisms, masks can reduce computational resources and contribute to the scalability of machine learning models. Notably, masked language modeling (Song et al., 2019; Wettig et al., 2023) and masked vision learning (Li et al., 2023) showcase the significance of masks.
@@ 4 Model Construction
Traditional graph learning methods struggle with extensive memory usage and slow computation. Node-wise sampling methods, including well-known ones like GraphSAGE (Hamilton et al., 2018), are not optimized for multi-hop scenarios due to the computational resources usage that grows exponentially with each additional hop— a phenomenon known as the "neighbor explosion." This makes them unsuitable for large-scale graphs where the capture of extended neighborhoods is crucial. Layer-wise sampling, such as in FastGCN(Chen et al., 2018), although designed to mitigate this explosion by sampling nodes per layer, often loses valuable connectivity information, leading to a sparse and inaccurate representation of the graph’s multi-hop structure. Furthermore, subgraph sampling methods such as Cluster-GCN (Chiang et al., 2019) focus on localized computations and overlook vital inter-subgraph connections, failing to capture the broader topology necessary for a comprehensive multi-hop domain analysis.
@@
GraphSnapShot is designed to solve the inefficacy for current graph masking methods on large graphs. Typical distributed graph processing System such as Marius (Mohoney et al., 2021), will re-sample all the local structure each time, and load corresponding node embeddings from disk, which is time-consuming. GraphSnapShot proposing a new method for quick local structure retrieval by integrating the benefits of both static and dynamic processing. In the preprocessing phase, a static graph snapshot is sampled and stored in memory from the disk. Then, in the computation phase, the algorithm re-samples the graph snapshot, using a mix of data partially from memory and partially from disk. Concurrently, a smaller, dynamic snapshot is sampled from the disk. This dynamic snapshot is crucial as it is used to swap with the portion of the graph stored in memory. This process ensures the computation can always learn "fresh" snapshot without overwhelming the memory capacity. GraphSnapShot allows the algorithm to continuously update and adapt to changes in the graph while maintaining a manageable memory footprint, making it an effective tool for processing large, dynamic graph data.
@@ 5.2 OTF
The "On The Fly" (OTF) method, updates portions of the cache dynamically while leaving other segments of the cache information unchanged. In detailed implementation of OTF, we design 4 different mode of it for real world application. They are (Full Fetch, Partial Fetch)x(Full Refresh, Partial Refresh).
@@ 5.5 FCR
The third method is FCR, which is short for "fully cache refreshing" in a batch. For FCR, in each batch, the the structure information are sampled from cache for computation, and between each batch, the cache is fully refreshed.
@@ 6 GraphSnapShot Overview
We use the Deep Graph Library (DGL(Wang et al., 2020)) framework to implement new samplers in GraphSnapShot, leveraging its efficient handling of graph data structures and operations. Our implementation involves loading the graph, applying the graph splitting based on node degree, and then processing each subgraph with the respective sampling technique. We evaluated the performance of our new models on datasets in ogbn-benchmark(Hu et al., 2021). The results show that GraphSnapShot can significantly reduce training time and memory usage without compromising the quality of GNN training.
@@ 7.1 Theoretial Analysis of Traditional Disk-Memory Strategy and Our Disk-Cache-Memory Strategy
In the traditional Graph System like Marius(Mohoney et al., 2021), a disk-cache model is deployed, where the graph structure is resampled each time from the disk into the memory for computation. In our current GraphSnapShot model, we deploy a disk-cache-memory model where K-V pair for reused graph structure are stored in the cache to promote the system performance.
@@ 7.6 Experimental Analysis of SSDReS Performance
Training Time and refresh rate: Our experiments indicated the existence of an optimal refresh rate value between 0 and 1, which minimizes training time, and with a little loss in acuracy. We call this point as "tradeoff" point between static and dynamic resampling. Over-dependence on dynamic resampling (low refresh rate) could increase the computational burden due to frequent updates required for rapidly changing graph structures, while excessive reliance on static resampling (high refresh rate) might cause the model to overly learn from limited portion of the graph structure in memory, deduce its ability to effectively learn from the entire graph structure.